# Engineering interaction potentials for stabilizing quantum quasicrystal phases

### v1

@@ II.1 Pair interaction potential
$8$-QC $2\cos\left(\pi/8\right)$ $2\cos\left(3\pi/8\right)$ $2\cos\left(2\pi/8\right)$
$10$-QC $2\cos\left(\pi/10\right)$ $2\cos\left(4\pi/10\right)$ $2\cos\left(2\pi/10\right)$
$12$-QC $2\cos\left(\pi/12\right)$ $2\cos\left(5\pi/12\right)$ $2\cos\left(3\pi/12\right)$
Table 2: Secondary wave vectors excited in $\hat{v}\left(k\right)$ for each quasicrystal structure considered. The patterns $n$-QC, with $n=8$, $10$ and $12$, correspond to the octagonal, decagonal, and dodecagonal structures, respectively.
To identify the key ingredients in this process, we considered pair interactions with a varying number of minima, i.e., considering cases in which some of the coefficients $d_{j}$’s are equal to zero. Once we select a particular minima structure positioned at $\{\tilde{k}_{j}\}$ in momentum space, the corresponding coefficients $\{d_{j}\}$ are obtained setting the values of the pair potential at the local minima $\{\hat{v}(\tilde{k}_{j})\}$. For all cases considered, we have $g=10$, $\sigma_{j}=0.1$ and $\hat{v}(\tilde{k}_{1})=-1$. This choice of dimensionless parameters is consistent with the range of values achievable in current experiments for the contact interaction and intensity of the laser, respectively Bonifacio et al. (2024a). For a systematic study, we determine the ground-state phase diagram of the different cases using as running parameters the dimensionless intensity of the pair potential $\lambda^{2}U\rho$ and the value of the pair potential at a local minimum $\hat{v}(\tilde{k}_{j})$, where $\tilde{k}_{j}$ represents a selected resonant wave vector of the corresponding structure (see Table 2).
@@ II.2 Superfluid fraction
To deepen our characterization of the quasicrystalline phases, we study the behavior of the superfluid fraction $\left(f_{s}\right)$ in these states. From the knowledge of the ground-state wave function $\psi\left(\bm{r}\right)$ the superfluid fraction can be estimated by means of the Legget criterion, which provides a lower bound for this quantity. Following well established literature Leggett (1970, 1998) we consider
@@
In Eq. (6), the prime symbol indicates that the coordinate system is rotated at an angle $\alpha$ with respect to the original coordinate system, consequently $\psi^{\prime}\left(\bm{r}^{\prime}\right)=\psi\left(x^{\prime}\cos\left(\alpha\right)-y^{\prime}\sin\left(\alpha\right),x^{\prime}\sin\left(\alpha\right)+y^{\prime}\cos\left(\alpha\right)\right)$. As occurs in the case of supersolids Blakie (2024), the angle $\alpha$ providing the highest value for $f_{s}$ coincides with the main directions of the structure, which in our case are given by the angles $\alpha_{i}=2\pi i/n$, with $i=0,...,n-1$, for an $n$-fold quasicrystalline phase. Hence, for calculation purposes, we can simply set $\alpha=0$.
The simultaneous characterization of the superfluid properties and the density pattern allow us to identify different kinds of phases: supersolid states, when superfluidity and a crystalline density pattern coexist, super quasicrystalline states, when the quasicrystal patterns present a finite superfluid fraction, and homogeneous superfluid phases. Moreover, all these phases have their insulating counterparts when the superfluid fraction is close enough to zero Gautier et al. (2021); Mendoza-Coto et al. (2022); Grossklags et al. (2024).
Figure 1: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), where the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary resonant scales are determined in Table 2. Phase diagram for the minima structure associated with the dodecagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The acronym HSF stands for homogeneous superfluid, whereas HS and $12$-QC stands for hexagonal solid and dodecagonal quasicrystal phases, respectively. The hexagonal states HSi refers to a hexagonal phase with characteristic wave vectors dominated by the minimum of $\hat{v}(k)$ at $\tilde{k}_{i}$.
@@ II.3 Gross-Pitaevskii equation
To verify the results computed within the spectral mean field variational method, we numerically solve the Gross-Pitaevskii Pitaevskii (1961); Gross (1963) equation (GPE) evolved in imaginary time, which provides the ground-state wave function and allows the calculation of the observables of our choice.
@@
[…]
@@ III Dodecagonal quasicrystal
To study the stabilization of the dodecagonal quasicrystalline pattern, we begin by considering an effective interaction potential with two length scales. The first minimum in $\hat{v}(k)$ is placed at $k=\tilde{k}_{1}=1$ while the second one is positioned at $k=\tilde{k}_{2}$, which corresponds to a length scale associated with a resonant wave vector of the dodecagonal structure (see Table 2). By fixing the depth of the first minimum to be $\hat{v}(\tilde{k}_{1})=-1$, it is possible to study the stabilization of the dodecagonal phase as the vertical position of the second minimum is varied. The form of $\hat{v}\left(k\right)$ is shown in Fig. 1(a) for three distinct values of the depth of the second minimum $\hat{v}(\tilde{k}_{2})$ and the corresponding ground-state phase diagram in the $\hat{v}(\tilde{k}_{2})$$\lambda^{2}U\rho$ plane is shown in Fig. 1(d). As can be observed, three different phases are identified, a homogeneous superfluid phase (HSF) at low values of $\lambda^{2}U\rho$, a dodecagonal quasicrystal ($12$-QC) phase for low values of $\hat{v}(\tilde{k}_{2})$ and a hexagonal solid phase (HS1) with characteristic momentum given by $\tilde{k}_{1}$. We notice that the presence of a second minimum at the first resonant wave vector of the dodecagonal structure is sufficient to stabilize such quasicrystalline phase when this minimum is deep enough. However, in the case of a degenerate minima structure, i.e. $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})$, the $12$-QC phase is only stabilized at large values of $\lambda^{2}U\rho$ in a regime where presumably particles will be largely localized, in agreement with previous works Mendoza-Coto et al. (2022). Moreover, we observe that as the value of the second minimum in $\hat{v}\left(k\right)$ is decreased, the transition from the HS1 phase to the $12$-QC phase moves to lower values of $\lambda^{2}U\rho$. This behavior is maintained up to the point at which we observe a direct transition from the superfluid to the $12$-QC phase.
@@
Figure 2: (a) Real space mapping for the dodecagonal structure in the phase diagram of Fig. 1(d) for parameter values $\hat{v}(\tilde{k}_{2})=-2.30$ and $\lambda^{2}U\rho=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the dodecagonal vectors. (d) Superfluid fraction for a fixed density value $\lambda^{2}U\rho=0.50$ in the dodecagonal quasicrystal for the ground state depicted in Fig. 1(d). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via the solution of the GPE.
@@
To further study the stabilization of the $12$-QC phase, we modify the interaction potential including a third characteristic length scale manifesting as a minimum in $\hat{v}\left(k\right)$ at a resonant wave vector different from $\tilde{k}_{2}$. Here we decided to consider a scenario in which the minima in $\hat{v}\left(k\right)$ at $k=\tilde{k}_{1}$ and $k=\tilde{k}_{2}$ are degenerate, since this is a quite common assumption for models developing quasicrystalline phases, and investigate the effects of varying the depth of the third minimum added to $\hat{v}\left(k\right)$. Firstly, we examine the situation in which the third minima is added at $k=\tilde{k}_{3}$ (see Table 2), examples of the resulting pair potential $\hat{v}\left(k\right)$ are shown in Fig. 1(b). The corresponding phase diagram in the $\hat{v}(\tilde{k}_{3})$$\lambda^{2}U\rho$ plane is presented in Fig. 1(e). In agreement with the phase diagram obtained for the model with two minima, we observe that the addition of a third minimum not too deep in the pair interaction potential does not produce the stabilization of the $12$-QC phase in the range of $\lambda^{2}U\rho$ considered. However, when this minimum becomes negative we observe the existence of an interval of values in which the $12$-QC phase is again stabilized. Finally, we observe that when the third minimum added is deep enough, a reentrant transition to a new hexagonal phase (HS3) with characteristic wave vector given by $\tilde{k}_{3}$ takes place. This behavior is already expected since an isolated dominant minimum is anticipated to stabilize a hexagonal ground state.
@@
Lastly, we consider the excitation of a third minimum at the remaining wave vector belonging to the first generation of resonant wave vectors of the dodecagonal quasicrystal, the one with $k=\tilde{k}_{4}$ (see Table 2). In this scenario, the interaction potential takes the form depicted in Fig. 1(c) and the corresponding ground-state phase diagram varying $\hat{v}(\tilde{k}_{4})$ is shown in Fig. 1(f). As can be observed, the addition of this third length scale is much more effective in promoting the stabilization of the quasicrystalline phase when compared to the scenario shown in Fig. 1(e). In this case, even the addition of a positive minimum with a relative high value already induces the stabilization of the $12$-QC for moderate values of $\lambda^{2}U\rho$. As expected, the strongest effect is achieved when the minimum is negative and in this case the preemptive hexagonal phase $\mathrm{HS}_{1}$ is not present and we have a direct transition from the HSF to the $12$-QC phase when $\lambda^{2}U\rho$ is increased. Moreover, as in the previous case, a further decrease of $\hat{v}(\tilde{k}_{4})$ eventually promotes a transition to the HS4 phase with characteristic wave vector close to $\tilde{k}_{4}$.
@@
The three scenarios discussed above show different routes for the stabilization of the $12$-QC phase in two dimensional bosonic gases. Although, in this case, the presence of two properly positioned minima is enough for such stabilization, differently from the classical case Barkan et al. (2014); Dotera (2011), it requires a second high-momentum minimum much deeper than the first one to stabilize the $12$-QC phase in the low $\lambda^{2}U\rho$ regime. Alternatively, the addition of a third minimum to $\hat{v}(k)$ at $\tilde{k}_{3}$ or $\tilde{k}_{4}$ enhance significantly the stability of the $12$-QC phase, even allowing us to observe this phase with a degenerate pair potential $(\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2}))$. In this sense, by comparing Fig. 1(e) and Fig. 1(f) we conclude that in order to enhance the stability of the $12$-QC phase, it is more effective to decrease $\hat{v}(\tilde{k}_{4})$ than $\hat{v}(\tilde{k}_{3})$.
@@
Regarding the structural properties of the $12$-QC phase, as an example, we consider the scenario presented in Fig. 1(a) with parameters $\hat{v}(\tilde{k}_{2})=-2.30$ and $\lambda^{2}U\rho=0.50$. In Fig. 2(a), we show the ground-state density pattern obtained after the energy minimization, and in Fig. 2(b) a graphic of the corresponding Fourier transform of the solution. The radii of the circles are proportional to the modulus of the Fourier amplitudes, excluding the circle corresponding to the $k=0$ Fourier mode, which is not presented. The red circumferences indicate the position of the minima of $\hat{v}(k)$ in this case. As can be observed, the presence of the minima in $\hat{v}(k)$ controls which are the main modes excited in the quasicrystalline density pattern, this kind of phenomenology was verified in all scenarios considered in this work.
@@
Figure 3: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), in which the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary resonant scales are determined in Table 2. Phase diagram for the minima structure associated with the decagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The notation employed to label the phases is analogous to the one established in Fig.1.
@@
To better understand the process of stabilization of the dodecagonal pattern and the greater stability of this structure in comparison with other quasicrystalline phases, as we will see later, we consider a simplified two-mode expansion including only the two main modes of the structure with $|\bm{k}_{j}|=\{\tilde{k}_{1},\tilde{k}_{2}\}$, as it is the case in Fig. 1(a). Our goal here is to assess the relative importance of the various interactions present in the energy per particle functional. Replacing Eq. (4) into Eq. (3) we can obtain the general expression:
@@
$\begin{split}\frac{E[\psi]}{N}=&\frac{1}{8Z}\sum\limits_{j}\bm{k}^{2}_{j}\lvert c\left(\bm{k}_{j}\right)\rvert^{2}+\frac{\lambda^{2}U\rho}{32Z^{2}}\sum\limits_{j,l,m}c\left(\bm{k}_{j}\right)c\left(\bm{k}_{j}+\bm{k}_{l}\right)^{\ast}\times\\ &\times c\left(\bm{k}_{m}\right)c\left(\bm{k}_{m}-\bm{k}_{l}\right)^{\ast}\hat{v}\left(\bm{k}_{l}\right),\end{split}$ (8)
where $Z=\frac{1}{4}\sum_{j}c\left(\bm{k}_{j}\right)^{2}$ and $c\left(0\right)=2$. This expression is general and naturally remains valid for the two mode ansatz under consideration. As we can observe from Eq. (8), the potential energy contribution is given as a sum of contractions of four Fourier amplitudes paired carrying momentum $\bm{k}_{l}$ and $-\bm{k}_{l}$. At this point, it is possible to identify which contractions produces the highest contributions to the potential energy and hence the leading mechanisms of stabilization of a given quasicrystalline phase. The most relevant type of contraction is naturally the one of the form $c\left(\bm{0}\right)c\left(\bm{k}_{1}\right)^{\ast}c\left(\bm{0}\right)c\left(-\bm{k}_{1}\right)^{\ast}\hat{v}\left(\bm{k}_{1}\right)$ and its equivalents. However, this kind of contraction is present in the energy per particle of all the solutions considered due to their pattern inversion symmetry. The next most relevant contribution, which is only present for solutions with symmetry against rotations of $\pi/3$, are of the type $c\left(\bm{0}\right)c\left(\bm{k}_{1}\right)^{\ast}c\left(\bm{k}^{\prime}_{1}\right)c\left(\bm{k}^{\prime}_{1}-\bm{k}_{1}\right)^{\ast}\hat{v}\left(\bm{k}_{1}\right)$, where the vectors $\bm{k}_{1}$ and $\bm{k}^{\prime}_{1}$ are of equal modulus $\left(\lvert\bm{k}_{1}\rvert=\lvert\bm{k}^{\prime}_{1}\rvert\right)$ and form an angle of $\pi/3$ between them. As we can see, this is a contribution combining one zero-momentum Fourier amplitude with three amplitudes corresponding to vectors in the same shell, see Fig. 2(c) for a pictorial representation of the corresponding contraction. In all the tests performed, this kind of term was responsible for around $40\%$ of the total energy of the reduced expansion for the dodecagonal pattern explaining the great stability of this phase in comparison to other quasicrystal solutions that do not contain such interaction terms.

### v2

@@ II.1 Pair interaction potential
88-QC $\lvert\bm{e}_{1}+\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(\pi/8\right)$ $\lvert\bm{e}_{1}-\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(3\pi/8\right)$ $\lvert\bm{e}_{1}+\bm{e}_{3}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(2\pi/8\right)$
1010-QC $\lvert\bm{e}_{1}+\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(\pi/10\right)$ $\lvert\bm{e}_{1}-\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(4\pi/10\right)$ $\lvert\bm{e}_{1}+\bm{e}_{3}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(2\pi/10\right)$
1212-QC $\lvert\bm{e}_{1}+\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(\pi/12\right)$ $\lvert\bm{e}_{1}-\bm{e}_{2}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(5\pi/12\right)$ $\lvert\bm{e}_{1}+\bm{e}_{4}\rvert/\lvert\bm{e}_{1}\rvert=2\cos\left(3\pi/12\right)$
Table 2: Secondary wave vectors moduli excited in $\hat{v}\left(k\right)$ for each quasicrystal structure considered. The patterns $n$-QC, with $n=8$, 1010 and $12$, correspond to the octagonal, decagonal, and dodecagonal structures, respectively. The vector basis $\{\bm{e}_{j}\}$ for each kind of structure is depicted in Tab. 1.
To identify the key ingredients in this process, we considered pair interactions with a varying number of minima, i.e., considering cases in which some of the coefficients $d_{j}$’s are equal to zero. Once we select a particular minima structure positioned at $\{\tilde{k}_{j}\}$ in momentum space, the corresponding coefficients $\{d_{j}\}$ are obtained setting the values of the pair potential at the local minima $\{\hat{v}(\tilde{k}_{j})\}$. For all cases considered, we have $g=10$, $\sigma_{j}=0.1$ and $\hat{v}(\tilde{k}_{1})=-1$. This choice of dimensionless parameters is consistent with the range of values achievable in current experiments for the contact interaction and intensity of the laser, respectively [bonifacio2024laser]. For a systematic study, we determine the ground-state phase diagram of the different cases using as running parameters the dimensionless intensity of the pair potential $\gamma$ and the value of the pair potential at a local minimum $\hat{v}(\tilde{k}_{j})$, where $\tilde{k}_{j}$ represents a selected characteristic wave vector of the corresponding structure (see Tab. 2). The particular set of momenta $\{\tilde{k}_{j}\}$ displayed in Tab. 2 corresponds to the modulus of specific higher order characteristic wave vectors of the corresponding quasicrystalline patterns. In this case, as shown in Tab. 2, we consider higher order momenta resulting from combining pairs of wave vectors of the basis (Tab. 1), for each quasicrystalline structure.
@@ II.2 Superfluid fraction
To deepen our characterization of the quasicrystalline phases, we study the behavior of the superfluid fraction $\left(f_{s}\right)$ in these states. From the knowledge of the ground-state wave function $\psi\left(\bm{r}\right)$ the superfluid fraction can be estimated by means of the Legget criterion, which provides a lower bound for this quantity. Following well established literature [leggett1970can, leggett1998superfluid] we consider
@@
In Eq. (6), the prime symbol indicates that the coordinate system is rotated at an angle $\alpha$ with respect to the original coordinate system, consequently $\psi^{\prime}\left(\bm{r}^{\prime}\right)=\psi\left(x^{\prime}\cos\left(\alpha\right)-y^{\prime}\sin\left(\alpha\right),x^{\prime}\sin\left(\alpha\right)+y^{\prime}\cos\left(\alpha\right)\right)$. As occurs in the case of supersolids [blakie2024superfluid], the angle $\alpha$ providing the highest value for $f_{s}$ coincides with the main directions of the structure, which in our case are given by the angles $\alpha_{i}=2\pi i/n$, with i=0,,n1i=0,...,n-1, for an $n$-fold quasicrystalline phase. Hence, for calculation purposes, we can simply set $\alpha=0$.
The simultaneous characterization of the superfluid properties and the density pattern allow us to identify different kinds of phases: supersolid states, when superfluidity and a crystalline density pattern coexist, super quasicrystalline states, when the quasicrystal patterns present a finite superfluid fraction, and homogeneous superfluid phases. Moreover, all these phases have their insulating counterparts when the superfluid fraction is close enough to zero [gautier2021strongly, mendoza2022exploring, grossklags2024self].
Figure 1: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), where the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary characteristic scales are presented in Tab. 2. Phase diagram for the minima structure associated with the dodecagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The acronym HSF stands for homogeneous superfluid, whereas HS and 1212-QC stands for hexagonal solid and dodecagonal quasicrystal phases, respectively. The hexagonal states HSi refers to a hexagonal phase with characteristic wave vectors dominated by the minimum of $\hat{v}(k)$ at $\tilde{k}_{i}$.
@@ II.3 Gross-Pitaevskii equation
To verify the results computed within the spectral mean field variational method, we numerically solve the Gross-Pitaevskii [pitaevskii1961vortex, gross1963hydrodynamics] equation (GPE) evolved in imaginary time, which provides the ground-state wave function and allows the calculation of the observables of our choice.
@@
Eq. (7) is then solved numerically using an adaptive Runge-Kutta method, with a simulation box of linear size $L=512\lambda$ and discretization $\Delta x=\Delta y=0.25\lambda$. The non-local term $v_{NL}\left(\bm{r}\right)=\int d\bm{r}^{\prime}v(\bm{r}-\bm{r}^{\prime})\lvert\psi\left(\bm{r}^{\prime},\tau\right)\rvert^{2}$ is computed efficiently with the use of fast-Fourier transformations since ${\cal F}\left[v_{NL}\right]\left(\bm{k}\right)={\cal F}\left[v\right]\left(\bm{k}\right){\cal F}\left[\lvert\psi\rvert^{2}\right]\left(\bm{k}\right)$.
@@ III Dodecagonal quasicrystal
[…]
@@
Figure 2: (a) Real space mapping for the dodecagonal structure in the phase diagram of Fig. 1(d) for parameter values $\hat{v}(\tilde{k}_{2})=-2.30$ and $\gamma=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the dodecagonal vectors. (d) Superfluid fraction for a fixed intensity of the pair interaction, $\gamma=0.50$, as a function of $\hat{v}(\tilde{k}_{2})$ for the ground state phase diagram in Fig. 1(d). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via the solution of the GPE.
@@
To further study the stabilization of the $12$-QC phase, we modify the interaction potential including a third characteristic length scale manifesting as a minimum in $\hat{v}(k)$ at a specific wave vector of the pattern different from $\tilde{k}_{2}$. Here we decided to consider a scenario in which the minima in $\hat{v}(k)$ at $k=\tilde{k}_{1}$ and $k=\tilde{k}_{2}$ are degenerate, since this is a quite common assumption for models developing quasicrystalline phases, and investigate the effects of varying the depth of the third minimum added to $\hat{v}(k)$. Firstly, we examine the situation in which the third minima is added at $k=\tilde{k}_{3}$ (see Tab. 2), examples of the resulting pair potential $\hat{v}(k)$ are shown in Fig. 1(b). The corresponding phase diagram in the $\hat{v}(\tilde{k}_{3})$$\gamma$ plane is presented in Fig. 1(e). In agreement with the phase diagram obtained for the model with two minima, we observe that the addition of a third minimum not too deep in the pair interaction potential does not produce the stabilization of the $12$-QC phase in the range of $\gamma$ considered. However, when this minimum becomes negative we observe the existence of an interval of values in which the $12$-QC phase is again stabilized. Finally, we observe that when the third minimum added is deep enough, a reentrant transition to a new hexagonal phase (HS3) with characteristic wave vector given by $\tilde{k}_{3}$ takes place. This behavior is already expected since an isolated dominant minimum is anticipated to stabilize a hexagonal ground state.
@@
Lastly, we consider the excitation of a third minimum at the remaining wave vector belonging to the first generation of characteristic wave vectors of the dodecagonal quasicrystal, the one with $k=\tilde{k}_{4}$ (see Tab. 2). In this scenario, the interaction potential takes the form depicted in Fig. 1(c) and the corresponding ground-state phase diagram varying $\hat{v}(\tilde{k}_{4})$ is shown in Fig. 1(f). As can be observed, the addition of this third length scale to $\hat{v}(k)$ is much more effective in promoting the stabilization of the quasicrystalline phase when compared to the scenario shown in Fig. 1(e). In this case, even the addition of a positive minimum with a relative high value already induces the stabilization of the $12$-QC for moderate values of $\gamma$. As expected, the strongest effect is achieved when the minimum is negative, in which case the preemptive hexagonal phase $\mathrm{HS}_{1}$ is not present and we have a direct transition from the HSF to the $12$-QC phase increasing $\gamma$. Moreover, as in the previous case, a further decrease of $\hat{v}(\tilde{k}_{4})$ eventually produce a transition to the HS4 phase with characteristic wave vector close to $\tilde{k}_{4}$. It is important to note that all transitions displayed in the phase diagrams in Fig. 1 are first-order. We verify this behavior by monitoring the Fourier amplitudes of the modulated phases at the homogeneous-modulated transition, which takes a significant value at the transition point. On the other hand, for the structural transitions between modulated phases, the symmetry incompatibility between quasicrystalline and the crystalline states allow us to rule out the possibility of a second-order phase transition. Moreover, we would like to mention that the scenario described above is not exclusive to the dodecagonal quasicrystal pattern, it is also observed in the case of the decagonal (Fig. 3) and octagonal (Fig. 5) structures.
At this point, it is interesting to notice how the monotony of the phase boundaries of the dodecagonal quasicrystalline phase varies from Fig. 1(e) to (f). In the first case, the leading Fourier mode of the $12$-QC phase varies between the low and high $\gamma$ regime. At low values of $\gamma$, when $\hat{v}(\tilde{k}_{3})$ is negative but small, the leading Fourier mode of the $12$-QC is the one with the lowest value of the characteristic momenta ($\tilde{k}_{3}$). In this regime, the kinetic energy contribution plays a more important role than the potential one, which is a $12$-QC solution with a lower characteristic momentum is energetically preferred. Moreover, at large values of $\gamma$ the potential energy contributions dominates over the kinetic energy producing a quasicrystal pattern whose leading Fourier amplitude corresponds to the wave vector of the absolute minimum of $\hat{v}(k)$, i.e., $k=\tilde{k}_{1}$. This discussion explains the occurrence of the $12$-QC to $\mathrm{HS}_{1}$ phase transition while increasing $\gamma$. Such phase transition is driven by the system’s tendency to decrease the potential energy, favoring a modulation whose characteristic wave vector coincides with the absolute minimum of $\hat{v}(k)$. The same mechanism also explains the transition from the $\mathrm{HS}_{3}$ phase to the $12$-QC phase at lower values of $\hat{v}(\tilde{k}_{3})$, as we increase $\gamma$. In the case of Fig. 1(f), the situation is different. Here, the minimum at $k=\tilde{k}_{4}$ excites a set of Fourier modes acting as higher-order harmonics of the $12$-QC structure, thereby enhancing its stability. This effects remains until the minimum at $k=\tilde{k}_{4}$ becomes the dominant one and the system naturally develops a transition to a structure primarily exciting this wave vector, the $\mathrm{HS}_{4}$ phase.
Finally, we would like to point out that, in the special case shown in Fig. 1(d), the ground-state phase diagram does not exhibit a reentrant transition to the $\mathrm{HS}_{2}$ phase, even for relatively low values of $\hat{v}(\tilde{k}_{2})$, to the best of our knowledge. This interesting behavior appears to arise from the fact that, in this configuration, the two main modes of the quasicrystal ground-state wave function, corresponding to $k=\tilde{k}_{1}$ and $k=\tilde{k}_{2}$, remain active even at low values of $\hat{v}(\tilde{k}_{2})$, as both types of modes contribute to the density Fourier components at $k=\tilde{k}_{2}$. Furthermore, the simultaneous presence of these two modes enables the formation of a quasicrystalline pattern with lower kinetic energy, ultimately leading to a $12$-QC phase with lower total energy than the $\mathrm{HS}_{2}$, even in the regime of low $\hat{v}(\tilde{k}_{2})$.
@@
The three scenarios discussed above show different routes for the stabilization of the $12$-QC phase in two dimensional bosonic gases. Although, in this case, the presence of two properly positioned minima is enough for such stabilization, differently from the classical case [barkan2014controlled, dotera2011quasicrystals], it requires a second high-momentum minimum much deeper than the first one to stabilize the $12$-QC phase in the low $\gamma$ regime. Alternatively, the addition of a third minimum to $\hat{v}(k)$ at $\tilde{k}_{3}$ or $\tilde{k}_{4}$ enhance significantly the stability of the $12$-QC phase, even allowing us to observe this phase with a degenerate pair potential $(\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2}))$. In this sense, by comparing Fig. 1(e) and Fig. 1(f) we conclude that in order to enhance the stability of the $12$-QC phase, it is more effective to decrease $\hat{v}(\tilde{k}_{4})$ than $\hat{v}(\tilde{k}_{3})$.
@@
Regarding the structural properties of the $12$-QC phase, as an example, we consider the scenario presented in Fig. 1(a) with parameters $\hat{v}(\tilde{k}_{2})=-2.30$ and $\gamma=0.50$. In Fig. 2(a), we show the ground-state density pattern obtained after the energy minimization, and in Fig. 2(b) a graphic of the corresponding Fourier transform of the solution. The radii of the circles are proportional to the modulus of the Fourier amplitudes, excluding the circle corresponding to the $k=0$ Fourier mode, which is not presented. The red circumferences indicate the position of the minima of $\hat{v}(k)$ in this case. As can be observed, the presence of the minima in $\hat{v}(k)$ controls which are the main modes excited in the quasicrystalline density pattern, this kind of phenomenology was verified in all scenarios considered in this work.
@@
Figure 3: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), in which the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary characteristic wave vectors are presented in Tab. 2. Phase diagram for the minima structure associated with the decagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The notation employed to label the phases is analogous to the one established in Fig.1.
@@
To better understand the process of stabilization of the dodecagonal pattern and the greater stability of this structure in comparison with other quasicrystalline phases, as we will see later, we consider a simplified two-mode expansion including only the two main modes of the structure with $|\bm{k}_{j}|=\{\tilde{k}_{1},\tilde{k}_{2}\}$, as it is the case in Fig. 1(a). Our goal here is to assess the relative importance of the various interactions present in the energy per particle functional. Replacing Eq. (4) into Eq. (3), we can obtain the general expression
@@
$\begin{split}\frac{E[\psi]}{N}=&\frac{1}{8Z}\sum\limits_{j}\bm{k}^{2}_{j}\lvert c\left(\bm{k}_{j}\right)\rvert^{2}+\frac{\gamma}{32Z^{2}}\sum\limits_{j,l,m}c\left(\bm{k}_{j}\right)c\left(\bm{k}_{j}+\bm{k}_{l}\right)^{\ast}\times\\ &\times c\left(\bm{k}_{m}\right)c\left(\bm{k}_{m}-\bm{k}_{l}\right)^{\ast}\hat{v}\left(\bm{k}_{l}\right),\end{split}$ (8)
[…]